For each given function value, find the remaining five trigonometric function values.
Tan (theta) = -2 and theta is in the second quadrant.

Answer by lwsshak3(11628)

For each given function value, find the remaining five trigonometric function values.
Tan (theta) = -2 and theta is in the second quadrant.
***
Hypotenuse of reference right triangle in quadrant 2=√((1^2)+(2^2))=√5
Tan (theta)=-2
cot (theta)=-1/2
sin (theta)=2/√5=2√5/5
cos (theta)=-1/√5=-√5/5
csc (theta)=√5/2
sec (theta)=-√5